OpenStudy (anonymous):

Write the expression as a single natural logarithm. 3 ln x - 2 ln c

OpenStudy (tkhunny):

Use this rule twice: \(a\cdot log(b) = log\left(b^{a}\right)\)

OpenStudy (anonymous):

Whats the value of a and B?

OpenStudy (tkhunny):

This is where you look at your problem statement and try to draw parallels with the example and rule. This answers your question directly.

OpenStudy (unklerhaukus):

and then use this rule \[\log p -\log q = \log(p/q)\]

OpenStudy (anonymous):

I still don't get where im suppose to plug in the numbers

OpenStudy (mosaic):

\(a*\ln(x) = \ln(x^a)\). Therefore, \(3*\ln(x) = \ln(x^3)\) \(2*\ln(c) = ?\)

OpenStudy (anonymous):

ln(x^2)?

OpenStudy (mosaic):

In the second part of the original problem, is it x or c?

OpenStudy (anonymous):

oh its 2*ln(c)=ln(c^2) ?

OpenStudy (mosaic):

okay. Subtract the two parts. Use the formula given by UnkleRhaukus to simplify.

OpenStudy (mosaic):

\( \Large 3*\ln x - 2*\ln c = \ln(x^3) - \ln(c^2) = \ln(\frac{x^3}{c^2})\)

OpenStudy (anonymous):

I was gonna do it! lol haha cx But ok thanks :)

OpenStudy (mosaic):

You are welcome.